OpenStudy (anonymous):

6x+5y=61 -5x+5y=-5 A. (8, 3) B. (6, 5) C. (6, 3) D. (8, 5)

OpenStudy (anonymous):

This is a system of equations. 6x + 5y = 61 -5x + 5y = -5 by property of the equation we can multiply both parts by any nonzero number. so we multiply second one by -1 6x + 5y = 61 5x - 5y = 5 then by the rule of summ we sum up our equations and eliminate y by doing this. 11x = 66 from here x = 66/11 = 6 5 * 6 - 5y = 5 -5y = 5 - 30 = -25 y = -25 / -5 = 5 check: 6 * 6 + 5 * 5 = 61 36 + 25 = 61 61 = 61 Therefore variant B (6 5))

OpenStudy (anonymous):

it is (6,5)